# Forecasting intraday foreign exchange volatility with functional GARCH approaches

### v2

@@ 4.1 Multi-level FPCA
As an extension of the functional mixed model (Morris and Carroll, 2006), the multi-level functional time series model has been successfully applied across several fields, improving both exploration and forecasting through incorporating cross-dependence variations (see, e.g., Cuesta-Albertos and Febrero-Bande, 2010; Shang, 2016; Shang and Kearney, 2022; Shang et al., 2022). Here, let us consider $d$-dimensional multivariate squared intraday return curves $\{y^{2}_{1,t}(u),y^{2}_{2,t}(u),\dots,y^{2}_{d,t}(u)\}$ ($d=3$ in our context), the multi-level functional data framework follows the idea of the common factor model by suggesting that these curves can be decomposed into an average functional mean $\mu_{(c)}(u)$, a forex-specific deviation from the averaged functional mean $\mu_{(j)}(u)$, a common trend across forex rates $U_{(c),t}(u)$, and a forex-specific residual trend $U_{(j),t}(u)$, for $1\leq j\leq d$. Thus, the $j$th squared OCIDR curve can be written as,
@@
Concentrating on the common and forex-specific variation components $\{U_{(c),t}(u)\}$ and $\{U_{(j),t}(u)\}$, we can estimate their sample long-run covariance operators $c_{U_{c},N}(u,v)$ and $c_{U_{j},N}(u,v)$, respectively. The linearly independent basis functions $\{\widehat{\psi}^{(c)}_{1}(u),\dots,\widehat{\psi}^{(c)}_{K^{c}}(u)\}$ and $\{\widehat{\psi}^{(j)}_{1}(u),\dots,\widehat{\psi}^{(j)}_{K^{j}}(u)\}$ associated with the corresponding eigenvalues can be obtained by eigen-decomposing the uncorrelated operators $c_{U_{c},N}(u,v)$ and $c_{U_{j},N}(u,v)$, respectively. As we explained in Section 2, in practice, the values of the dimensions $K^{c}$ and $K^{j}$ can be chosen via the eigenvalue ratio method of Li et al. (2020).
@@
$\begin{split}&\omega^{(j)}=\sum_{l_{1}=1}^{K^{c}}d_{l_{1}}\psi^{(c)}_{l_{1}}+\sum_{l_{2}=1}^{K^{j}}d_{l_{2}}\psi^{(j)}_{l_{2}},\\ &\boldsymbol{\alpha}^{(j)}=\sum_{l_{1},k_{1}=1}^{K^{c}}a_{l_{1},k_{1}}\psi^{(c)}_{l_{1}}\psi^{(c)}_{k_{1}}+\sum_{l_{2},k_{2}=1}^{K^{j}}a_{l_{2},k_{2}}\psi^{(j)}_{l_{2}}\psi^{(j)}_{k_{2}}\\ &\boldsymbol{\beta}^{(j)}=\sum_{l_{1},k_{1}=1}^{K^{c}}b_{l_{1},k_{1}}\psi^{(c)}_{l_{1}}\psi^{(c)}_{k_{1}}+\sum_{l_{2},k_{2}=1}^{K^{j}}b_{l_{2},k_{2}}\psi^{(j)}_{l_{2}}\psi^{(j)}_{k_{2}},\\ &\boldsymbol{\gamma}^{(j)}=\sum_{l_{1},k_{1}=1}^{K^{c}}g_{l_{1},k_{1}}\psi^{(c)}_{l_{1}}\psi^{(c)}_{k_{1}}+\sum_{l_{2},k_{2}=1}^{K^{j}}g_{l_{2},k_{2}}\psi^{(j)}_{l_{2}}\psi^{(j)}_{k_{2}}.\end{split}$ (4.2)
@@
4.2 Long-range dependent FPCA
Treating $\{y_{t}^{2}{(u)}\}$ as short-range dependent curves can sometimes be inappropriate, as we learned in Table 3.1 that intraday return curves are long-range conditional heteroscedastic. In this section, we introduce a long-range dependent FPCA (LFPCA) method to obtain embedded basis functions with a long-range dependence structure.
Inspired by Li et al. (2020), we write $\{y_{t}^{2}(u)\}$ as a functional linear process,
@@ 4.2 Long-range dependent FPCA
where $\{b_{j},j\geq 0\}$ is a coefficient sequence and $\eta_{t-j}(u)\in\mathcal{L}^{2}[0,1]$ represents a series of independent identically distributed random functions. Under stationarity, we say that $\{y_{t}^{2}(u)\}$ is short-range dependent if $\sum_{j=0}^{\infty}|b_{j}|<\infty$; and $\{y_{t}^{2}(u)\}$ is long-range dependent if the summation is unbounded. We denote the memory parameter $a$, which describes the dependence strength of $\{y_{t}^{(2)}\}$ by reflecting $b_{j}\sim j^{a-1}$ in (4.3). Specifically, $\{y_{t}^{(2)}\}$ is deemed to be a stationary long-range dependent curve sequence when $11/2\leq a<$ 1; and it becomes non-stationary when $a\geq 1$.
In addition, we empirically test the stationarity of $\{y_{t}^{2}(u)\}$ across the multiple forecasting training samples that we employ in Section 5. Figure 4.1 shows the p-values of stationary tests implemented for squared OCIDR curves on 19 training samples.333 The periods covered by 19 training samples, which are updated every two months: 07-January-2014 to 17-May-2017; 27-March-2014 to 21-July-2017; 13-June-2014 to 22-September-2017; 28-August-2014 to 24-November-2017; 18-November-2014 to 31-January-2018; 10-February-2015 to 04-April-2018; 07-May-2015 to 06-June-2018; 24-July-2015 to 08-August-2018; 20-October-2015 to 10-October-2018; 05-January-2016 to 12-December-2018; 15-March-2016 to 26-February-2019; 19-May-2016 to 30-April-2019; 22-July-2016 to 02-July-2019; 23-September-2016 to 03-September-2019; 29-November-2016 to 05-November-2019; 01-February-2017 to 22-January-2020; 05-April-2017 to 25-March-2020; 13-June-2017 to 27-May-2020, 15-August-2017 to 11-August-2020. The results indicate that $\{y_{t}^{2}(u)\}$ could be stationary or non-stationary, such that the $\{y_{t}^{2}(u)\}$ samples of USD/GBP are stationary, while the other two currency pairs exhibit non-stationarity. In the following, we detail the LFPCA approaches to derive data-driven bases under these two scenarios.
Figure 4.1: $P$-values of stationarity tests for the squared Overnight Cumulative Intraday Return (OCIDR) curve, $\{y_{t}^{2}(u)\}$, across the various training samples utilised in the forecasting exercise in Section 5. $TSample\_x$ indicates the results for each training sample with $x$ ranging from 1 to 19. For example, the first training sample $TSample\_1$ covers the period 07-January-2014 to 17-May-2017 with the periods covered by the other training samples given in footnote 3. P-values less than the outlined 5% threshold indicate non-stationarity for that training sample.
4.2.1 Stationary long-range dependent curves
Following Li et al. (2020), the curve time series are stationary long-range dependent if the kernel coefficient operators in (4.3) are not square-summable and $11/2\leq a<$ 1. Let the unnormalised long-run covariance function be:
@@ 4.2.1 Stationary long-range dependent curves
According to Assumption 2 and the large sample properties in Li et al. (2020), we learn that the long-range dependent $y_{t}^{2}(u)$ can be decomposed into two parts: 1) the projection of $y_{t}^{2}(u)$ onto a finite number of sub-spaces that capture the long-range dependence structure of $y_{t}^{2}(u)$; and 2) the projection onto the remaining infinite-dimensional sub-spaces stores the short-range dependence. Thus, the long-range dependence of $y_{t}^{2}(u)$ can be captured by projecting onto a dominant sub-space $\mathcal{S}_{1}$ spanned over $\{\psi_{1},\psi_{2},\dots,\psi_{K_{1}}\}$. Then, dimension reduction can be achieved via the FPCA method on the normalised long-run covariance:
@@
where $a_{1}$ is a memory parameter measuring the dependence strength of the signal for the projection of $y_{t}^{2}(u)$ on $\mathcal{S}_{1}$. The normalised long-run covariance $\widetilde{c}(u,v)_{N}$ is not observable, so it has to be estimated through,
@@
The bandwidth $h$ is chosen using the plug-in algorithm proposed in Rice and Shang (2017) to minimise the estimator’s asymptotic mean-squared normed error. One can use the eigenvalue ratio method to determine the dimension $K_{1}$ of the dominant sub-space. Li et al. (2021) propose a semi-parametric feasible local Whittle estimator $\widehat{a}_{1}$ for the memory parameter in the context of functional time series. We suppress the technical details to save some space, and interested readers can refer to Li et al. (2021). Consequently, we obtain the consistent estimator of $\widetilde{c}_{h}(u,v)_{N}$ by substituting $\widehat{a}_{1}$ into (4.5), with the long-range dependent basis functions $\widehat{\psi}_{l}$ accessed through the eigen-decomposition method.
4.2.2 Non-stationary long-range dependent curves
We now discuss the case where $\{y_{t}^{2}(u)\}$ constitutes a non-stationary curve sequence (when $a>1/2$). Chang et al. (2016) is the seminal work in this field, paving a foundation for the subsequent developments in cointegration (Beare et al., 2017; Nielsen et al., 2023; Seo and Shang, 2024). They decompose a non-stationary functional time series from infinite-dimensional Hilbert space into a finite-dimensional I(1) subspace and infinite stationarity subspace. The main focus of the literature then turns to determine the finite dimension of the non-stationarity subspace (see, e.g., Beare et al., 2017; Franchi and Paruolo, 2020; Li et al., 2023). In this paper, we apply the approach proposed in Li et al. (2023), which assumes the data of interest follows a non-stationary fractionally integrated functional process as specified in model (4.3). This is a preferred method as it allows an unknown order of integration, which can be estimated using a feasible local Whittle estimator as in Section 4.2.1. Furthermore, similar to the stationary long-range dependent case, the linear operator $b_{j}$ is not square-summable. However, the memory parameter can be greater than one under non-stationarity. Li et al. (2023) show that such non-stationary functional curves can be divided into a finite dominant subspace $\overline{\mathcal{S}}_{1}$ that retains the non-stationary sample information and the remaining infinite asymptotically negligible subspaces. According to Theorem 3.2 in Li et al. (2023), we can define the normalised sample covariance operator on $\overline{\mathcal{S}}_{1}$,
@@ 4.2.2 Non-stationary long-range dependent curves
where $\otimes$ denotes the tensor product. For the estimation of the memory parameter $a_{1}$, we refer to the local Whittle estimator in Li et al. (2023). By eigen-decomposing $\overline{c}_{N}(u,v)$, we obtain $\widehat{\psi}_{l}$ eigenfunctions, for $1\leq l\leq\overline{K}_{1}$ , the span of which can generate the dominant space $\overline{\mathcal{S}}_{1}$. Similar to Section 4.2.1, $\overline{K}_{1}$ can be determined using the eigenvalue ratio method.
@@ 5 Predictive accuracy
We compare the forecasting models by assessing the accuracy of intraday and inter-daily volatility forecasts. The latter is feasible because the closing points of the intraday volatility forecasts for the OCIDR curves form a sequence of conditional volatility forecasts for daily returns, as we emphasised in Section 3. In order to obtain the forecasting errors of model candidates, ideally, we need unbiased and consistent estimators of latent volatility. This is feasible when we assess the inter-daily volatility forecasts, as the realised volatility $\text{RV}_{t}$ is an unbiased and consistent estimator that has been widely used in the literature (Barndorff-Nielsen and Shephard, 2002; Hansen and Lunde, 2006a; Zhang et al., 2024). Incorporating the overnight effect, we can calculate the realised volatility by summing squared intraday log returns with the squared overnight return,
@@
where $\Delta$ denotes an intraday sampling frequency. However, there is still no consistent and unbiased estimator for intraday volatility. We then consider the actual values of intraday volatility using the proxies of the squared OCIDR curves. We are aware that substituting the unknown consistent estimator of intraday conditional variance with squared OCIDR curves may result in unreliable and inferior results, similar to the same issue for scalar GARCH-type models discussed by Hansen and Lunde (2006a) and for multivariate cases discussed by Patton and Sheppard (2009).
Nonetheless, it is tractable to evaluate forecasting performance with noisy proxies. Patton (2011) documented that a noisy proxy can still be used to evaluate the volatility forecasts if a “robust” loss function is considered, which maintains an invariant property of ranking forecasts using an imperfect volatility proxy. Following Patton (2011), we apply two robust loss functions, namely mean squared forecasting error (MSFE) and QLIKE. Unlike MSFE, QLIKE is asymmetric; it penalises large under-forecasts much heavier than over-forecasts. Below, we define MSFE and QLIKE loss functions. For inter-daily conditional volatility forecasts, we have
@@
$\begin{split}\mbox{MSFE: }&L[\widetilde{\sigma}_{t}^{2}(u),\widehat{\sigma}_{t}^{2}(u)]=\frac{1}{T}\sum_{t=1}^{T}\left\{\frac{1}{J}\sum_{j=1}^{J}\left[\widetilde{\sigma}_{t}^{2}(u_{j})-\widehat{\sigma}_{t}^{2}(u_{j})\right]^{2}\right\},\\ \mbox{QLIKE: }&L[\widetilde{\sigma}_{t}^{2}(u),\widehat{\sigma}_{t}^{2}(u)]=\frac{1}{T}\sum_{t=1}^{T}\left\{\frac{1}{J}\sum_{j=1}^{J}\left[\log\widehat{\sigma}^{2}_{t}(u_{j})+\frac{\widetilde{\sigma}^{2}_{t}(u_{j})}{\widehat{\sigma}^{2}_{t}(u_{j})}\right]\right\},\\ \end{split}$ (5.2)
where the operations are performed in a pointwise manner, that is the proxy of squared OCIDR curves $\widetilde{\sigma}_{t}^{2}(u)$ and the predicted volatility curves $\widehat{\sigma}^{2}_{t}(u)$ are measured on $J$ grid points of $\widetilde{\sigma}_{t}^{2}(u_{j})$ and $\widehat{\sigma}^{2}_{t}(u_{j})$.

### v3

@@ 3.2.1 Multi-level FPCA
As an extension of the functional mixed model (Morris and Carroll,, 2006), the multi-level functional time series model has been successfully applied across several fields, improving both exploration and forecasting through incorporating cross-dependence variations (see, e.g., Cuesta-Albertos and Febrero-Bande,, 2010; Shang,, 2016; Shang and Kearney,, 2022; Shang et al.,, 2022). Here, let us consider $d$-dimensional multivariate squared intraday return curves $\{y^{2}_{1,t}(u),y^{2}_{2,t}(u),\dots,y^{2}_{d,t}(u)\}$ ($d=3$ in our context), the multi-level functional data framework follows the idea of the common factor model by suggesting that these curves can be decomposed into an average functional mean $\mu_{(c)}(u)$, a forex-specific deviation from the averaged functional mean $\mu_{(j)}(u)$, a common trend across forex rates $U_{(c),t}(u)$, and a forex-specific residual trend $U_{(j),t}(u)$, for $1\leq j\leq d$. Thus, the $j$th squared OCIDR curve can be written as,
@@
Concentrating on the common and forex-specific variation components $\{U_{(c),t}(u)\}$ and $\{U_{(j),t}(u)\}$, we can estimate their long-run covariance operators $c_{U_{c},N}(u,v)$ and $c_{U_{j},N}(u,v)$, respectively. Linearly independent basis functions $\{\widehat{\psi}^{(c)}_{1}(u),\dots,\widehat{\psi}^{(c)}_{K^{c}}(u)\}$ and $\{\widehat{\psi}^{(j)}_{1}(u),\dots,\widehat{\psi}^{(j)}_{K^{j}}(u)\}$ associated with the corresponding eigenvalues can be obtained by eigen-decomposing the uncorrelated operators $c_{U_{c},N}(u,v)$ and $c_{U_{j},N}(u,v)$, respectively. As we explained in Section 3.1, in practice, the values of dimensions $K^{c}$ and $K^{j}$ can be chosen using the eigenvalue ratio method of Li et al., (2020).
@@
[…]
@@
3.2.2 Long-range dependent FPCA
Treating $\{y_{t}^{2}{(u)}\}$ as short-range dependent curves can sometimes be inappropriate, as we learnt in Table 2.1 that intraday return curves are long-range conditional heteroscedastic. In this section, we introduce a long-range dependent FPCA (LFPCA) method to obtain embedded basis functions with a long-range dependence structure.
Inspired by Li et al., (2020), we write $\{y_{t}^{2}(u)\}$ as a functional linear process,
@@ 3.2.2 Long-range dependent FPCA
where $\{b_{j},j\geq 0\}$ is a sequence of coefficients and $\eta_{t-j}(u)\in\mathcal{L}^{2}[0,1]$ represents a series of independent identically distributed random functions. Under stationarity, we say that $\{y_{t}^{2}(u)\}$ is short-range dependent if $\sum_{j=0}^{\infty}|b_{j}|<\infty$; and $\{y_{t}^{2}(u)\}$ is long-range dependent if the summation is unbounded. We denote the memory parameter $a$, which describes the dependence strength of $\{y_{t}^{(2)}\}$ by reflecting $b_{j}\sim j^{a-1}$ in (3.9). Specifically, $\{y_{t}^{(2)}\}$ is deemed to be a stationary long-range dependent curve sequence when $1/2\leq a<1$; and it becomes non-stationary when $a\geq 1$.
In addition, we empirically test the stationarity of $\{y_{t}^{2}(u)\}$ across the multiple forecasting training samples that we employ in Section 4. Figure 3.1 shows the p-values of stationary tests implemented for the squared OCIDR curves in 19 training samples. [^6] The periods covered by 19 training samples, which are updated every two months: 07-January-2014 to 17-May-2017; 27-March-2014 to 21-July-2017; 13-June-2014 to 22-September-2017; 28-August-2014 to 24-November-2017; 18-November-2014 to 31-January-2018; 10-February-2015 to 04-April-2018; 07-May-2015 to 06-June-2018; 24-July-2015 to 08-August-2018; 20-October-2015 to 10-October-2018; 05-January-2016 to 12-December-2018; 15-March-2016 to 26-February-2019; 19-May-2016 to 30-April-2019; 22-July-2016 to 02-July-2019; 23-September-2016 to 03-September-2019; 29-November-2016 to 05-November-2019; 01-February-2017 to 22-January-2020; 05-April-2017 to 25-March-2020; 13-June-2017 to 27-May-2020, 15-August-2017 to 11-August-2020. The results indicate that $\{y_{t}^{2}(u)\}$ could be stationary or non-stationary, such that the $\{y_{t}^{2}(u)\}$ samples of USD/GBP are stationary, while the other two currency pairs exhibit non-stationarity. In the following, we detail the LFPCA approaches to derive data-driven bases under these two scenarios.
Figure 3.1: $P$-values of stationarity tests for the squared Overnight Cumulative Intraday Return (OCIDR) curve, $\{y_{t}^{2}(u)\}$, across the various training samples utilised in the forecasting exercise in Section 4. TSample_xTSample\_x indicates the results for each training sample with $x$ ranging from 1 to 19. For example, the first training sample TSample_1TSample\_1 covers the period 07-January-2014 to 17-May-2017 with the periods covered by the other training samples given in footnote 6. P-values less than the outlined 5% threshold indicate non-stationarity for that training sample.
Stationary long-range dependent curves
Following Li et al., (2020), the curve time series are stationary long-range dependent if the kernel coefficient operators in (3.9) are not square-summable and $1/2\leq a<1$. Let the unnormalised long-run covariance function be:
@@
According to Assumption 2 and the large-sample properties in Li et al., (2020), we learn that the long-range dependent $y_{t}^{2}(u)$ can be decomposed into two parts: 1) the projection of $y_{t}^{2}(u)$ onto a finite number of sub-spaces that capture the long-range dependence structure of $y_{t}^{2}(u)$; and 2) the projection onto the remaining infinite-dimensional sub-spaces stores the short-range dependence. Thus, the long-range dependence of $y_{t}^{2}(u)$ can be captured by projecting onto a dominant sub-space $\mathcal{S}_{1}$ spanned over $\{\psi_{1},\psi_{2},\dots,\psi_{K_{1}}\}$. Then, dimension reduction can be achieved via the FPCA method on the normalised long-run covariance:
@@
where $a_{1}$ is a memory parameter that measures the dependence strength of the signal for the projection of $y_{t}^{2}(u)$ on $\mathcal{S}_{1}$. The normalised long-run covariance $\widetilde{c}(u,v)_{N}$ is not observable, so it has to be estimated through,
@@
The bandwidth $h$ is chosen using the plug-in algorithm proposed in Rice and Shang, (2017) to minimise the estimator’s asymptotic mean-squared normed error. One can use the eigenvalue ratio method to determine the dimension $K_{1}$ of the dominant sub-space. Li et al., (2021) propose a semi-parametric feasible local Whittle estimator $\widehat{a}_{1}$ for the memory parameter in the context of functional time series. We suppress the technical details to save some space, and interested readers can refer to Li et al., (2021). Consequently, we obtain the consistent estimator of $\widetilde{c}_{h}(u,v)_{N}$ by substituting $\widehat{a}_{1}$ into (3.11), with long-range dependent basis functions $\widehat{\psi}_{l}$ accessed through the eigen-decomposition method.
Non-stationary long-range dependent curves
We now discuss the case where $\{y_{t}^{2}(u)\}$ constitutes a non-stationary curve sequence (when $a>1/2$). Chang et al., (2016) is the seminal work in this field, paving a foundation for the subsequent developments in cointegration (Nielsen et al.,, 2023; Seo and Shang,, 2024). They decompose a non-stationary functional time series from infinite-dimensional Hilbert space into a finite-dimensional I(1) subspace and infinite stationarity subspace. The main focus of the literature then turns to determine the finite dimension of the non-stationarity subspace (see, e.g., Li et al.,, 2023). In this paper, we apply the approach proposed in Li et al., (2023), which assumes the data of interest follow a non-stationary fractionally integrated functional process as specified in model (3.9). This is a preferred method as it allows an unknown order of integration, which can be estimated using a feasible local Whittle estimator as in the stationary case. Furthermore, similar to the stationary long-range dependent case, the linear operator $b_{j}$ is not square-summable. However, the memory parameter can be greater than one under non-stationarity. Li et al., (2023) show that such non-stationary functional curves can be divided into a finite dominant subspace $\overline{\mathcal{S}}_{1}$ that retains the non-stationary sample information and the remaining infinite asymptotically negligible subspaces. According to Theorem 3.2 in Li et al., (2023), we can define the normalised sample covariance operator on $\overline{\mathcal{S}}_{1}$,
@@
where $\otimes$ denotes the tensor product. For the estimation of the memory parameter $a_{1}$, we refer to the local Whittle estimator in Li et al., (2023). By eigen-decomposing $\overline{c}_{N}(u,v)$, we obtain $\widehat{\psi}_{l}$ eigenfunctions, for $1\leq l\leq\overline{K}_{1}$ , the span of which can generate the dominant space $\overline{\mathcal{S}}_{1}$. Similar to the stationary case, $\overline{K}_{1}$ can be determined using the eigenvalue ratio method.
@@ 4 Predictive accuracy
We compare the forecasting models by assessing the accuracy of intraday and inter-daily volatility forecasts. The latter is feasible because the closing points of the intraday volatility forecasts for the OCIDR curves form a sequence of conditional volatility forecasts for daily returns, as we emphasised in Section 2. In order to obtain the forecasting errors of the model candidates, ideally, we need unbiased and consistent estimators of latent volatility. This is feasible when we assess the inter-daily volatility forecasts, as the realised volatility $\text{RV}_{t}$ is an unbiased and consistent estimator that has been widely used in the literature (Barndorff-Nielsen and Shephard,, 2002; Hansen and Lunde, 2006a, ). Incorporating the overnight effect, we can calculate the realised volatility by summing squared intraday log returns with the squared overnight return,
@@
where $\Delta$ denotes an intraday sampling frequency. However, there is still no consistent and unbiased estimator for intraday volatility. We then consider the actual values of intraday volatility using the proxies of the squared OCIDR curves. We are aware that substituting the unknown consistent estimator of intraday conditional variance with squared OCIDR curves may result in unreliable and inferior results, similar to the same issue for scalar GARCH-type models discussed by Hansen and Lunde, 2006a and for multivariate cases discussed by Patton and Sheppard, (2009).
Nonetheless, it is tractable to evaluate forecasting performance with noisy proxies. Patton, (2011) documented that a noisy proxy can still be used to evaluate the volatility forecasts if a “robust” loss function is considered, which maintains an invariant property of ranking forecasts using an imperfect volatility proxy. Following Patton, (2011), we apply two robust loss functions, namely mean squared forecasting error (MSFE) and QLIKE. Unlike MSFE, QLIKE is asymmetric; it penalises large under-forecasts much heavier than over-forecasts. Below, we define MSFE and QLIKE loss functions. For inter-daily conditional volatility forecasts, we have
@@
[…]
where operations are performed in a pointwise manner, that is, the proxy of the squared OCIDR curves $\widetilde{\sigma}_{t}^{2}(u)$ and the predicted volatility curves $\widehat{\sigma}^{2}_{t}(u)$ are measured on the $J$ grid points of $\widetilde{\sigma}_{t}^{2}(u_{j})$ and $\widehat{\sigma}^{2}_{t}(u_{j})$.